OpenStudy (diana.xl):

A baseball player hits a home run. Which type of function would best represent the height of the ball over time? linear function quadratic function exponential function both a. and b.

OpenStudy (koikkara):

When a baseball is hit or thrown,Does it go straight alwayz ? @Diana.xL or Just goes like you see in the picture ? 

OpenStudy (koikkara):

i would go with exponential function

Directrix (directrix):

@Koikkara Nice diagram. I don't know how to embed a diagram with that drawing tool. I looked at this problem last night and was baffled by linear function because of the term "line drive" or "grounder" in baseball. No, not exponential.

Directrix (directrix):

That drawing is of a parabola and it is a quadratic function.

Directrix (directrix):

I am thinking that linear function could not be correct. The ball would never stop.

OpenStudy (baru):

@Koikkara i want to know too, how did u embed the diagram

Directrix (directrix):

I have a medal for you if you tell me. @Koikkara

Directrix (directrix):

@baru has a medal for you, too, I bet.

OpenStudy (baru):

i sure do xD

OpenStudy (koikkara):

lol, using OpenStudy Enhancement - extension available for google chrome... And no need of medal as somepoints help me earn more than a medal. You can get extension from user @Jayanator , he designed it.

Directrix (directrix):

I am not a Chrome user so I'm left out in the cold.

Directrix (directrix):

What is this: somepoints @Koikkara

OpenStudy (koikkara):

Ask jay he may have all sort of extensions. After installing extensions. Check the image... n thanks. It was a great pleasure to meet you guys !

OpenStudy (koikkara):

along with Drawing tool> Image will be added to it.

OpenStudy (baru):

thanks :) nice to meet you guys too

ganeshie8 (ganeshie8):

Near earth, if we assume the acceleration in vertical direction is constant : \(a(t)=-g\) \(v(t) = -gt+v(0)\) \(h(t)=-\frac{1}{2}gt^2+v(0)t+h(0)\) The degree of \(h(t)\) is \(2\), so it is a quadratic function...

Directrix (directrix):

That is what I said earlier in this thread. Nobody listened. @ganeshie8

OpenStudy (skullpatrol):

I listened pal :-)

Directrix (directrix):

>I listened pal :-) Thanks so much. I'm no longer a lone bark. If the answer is said but not heard, was there ever an answer?

OpenStudy (baru):

its definitely quadratic, but ganeshie, you have shown its quadratic with respect to 'time' the standard text book way is to assume initial velocty vector split it into componants and apply gravity constant to the y component. you will get a quadratic y=f(x)

ganeshie8 (ganeshie8):

right, we get a quadratic for the trajectory too as the velocity in horizontal direction is constant

OpenStudy (skullpatrol):

@Directrix teaching to deaf ears is not fun :(

OpenStudy (baru):

yep, constant horizontal velocity :)